# Boosting Transformer’s Robustness in PPG Signals Artifact Detection with Self-Supervised Learning

### v1

@@ III-C1 Masking
Then, the SSL framework for data reconstruction from the masked data utilizes the masked data generated by the previous pseudo-algorithm. This framework aims to train a model to predict the original data from its masked version, thereby learning a robust representation of the data. Through this procedure Algorithm 2, the model learns to predict the missing values in the masked data by reconstructing the original data. As the training progresses, the model becomes better at understanding the underlying patterns and structures in the data, even when some of the information is obscured. This SSL training paradigm effectively leverages the data as the supervisory signal, bypassing the need for external labels and allowing the model to learn unsupervised. This is particularly powerful for utilizing large unlabeled datasets to train models for tasks where labeled data is scarce to obtain.
III-C2 Contrastive Learning
@@
Contrastive learning is a SSL technique that teaches a model to distinguish between similar and dissimilar data points [30]. By doing so, the model learns rich, discriminative representations of data without the need for explicit labels. The essence of contrastive learning lies in its loss functions Table I, which drive the model to minimize the distance between positive pairs (similar items) and maximize the distance between negative pairs (dissimilar items). Here are three contrastive loss functions that are commonly used. Let’s denote $L_{ij}$ represents the loss for a positive pair of examples $(i,j)$. $\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)$ is the exponential function applied to the similarity score between the encoded representations $z_{i}$ and $z_{j}$. $\tau$ is the temperature parameter that scales the similarity score and controls the separation of the distribution of positive and negative examples.
  1. 1.
@@
    Normalized Temperature-scaled Cross Entropy (NT-Xent) Loss [31]: This loss function is central to many contrastive learning algorithms and has been popularized by its use in SimCLR. It normalizes feature vectors and scales the dot product between them with a temperature parameter, encouraging the model to identify positive pairs among a set of negative samples.
  2. 2.
@@
    Information Noise-Contrastive Estimation (InfoNCE) Loss [32, 33]: InfoNCE is a variant that focuses on distinguishing a positive pair from a set of negatives. It’s designed to learn energy-based models and has been effectively used in representation learning, improving the quality of learned representations.
  3. 3.
@@
    Swapped Cross Entropy (SwCE) Loss [34]: Introduced in SwAV, this function enforces consistency between cluster assignments produced by a pair of encoded views of the same image. It uses a swapping mechanism that leads to more robust representations by leveraging the invariance in the data.
These loss functions are integral to the SSL framework, enabling models to effectively leverage large amounts of unlabeled data for learning useful data representations. Each loss function has its advantages and can be chosen based on the specific requirements of the task at hand. Within this framework, we introduce our proposed Smooth InfoNCE Loss, an innovative take on the standard InfoNCE loss. The Smooth InfoNCE Loss incorporates a smoothing factor that modulates the influence of negative samples. This modification is designed to prevent the model from becoming overly confident in its negative sample predictions, thereby enhancing the generality and robustness of the features it learns. By doing so, we aim to mitigate the risk of overfitting negative samples within the training dataset.
TABLE I: Derivation of contrastive losses for contrastive self-supervised learning
Normalized Temperature-scaled Cross Entropy (NT-Xent) loss $L_{ij}=-\log\left(\frac{\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)}{\sum^{2N}_{k=1}1_{[k\neq i]}\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{k}\right)/\tau\right)}\right)$ (1) Information Noise-Contrastive Estimation (InfoNCE) loss $L_{ij}=-\log\left(\frac{\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)}{\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)+\sum^{N}_{k=1}\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{k}^{neg}\right)/\tau\right)}\right)$ (2) Swapped Cross Entropy (SwCE) loss $L_{ij}=-\log\left(\frac{\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)}{\sum^{N}_{k=1}\exp\left(\text{sim}\left(\mathbf{z}_{j},\mathbf{z}_{k}\right)/\tau\right)}\right)$ (3) Smooth InfoNCE loss - Our proposed loss function $L_{ij}=-\log\left(\frac{\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)}{\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)+\lambda\sum^{N}_{k=1,k\neq i}\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{k}\right)/\tau\right)}\right)$ (4)
Our Smooth InfoNCE Loss can be seen as a special case of the standard InfoNCE loss. Both losses share a commonality in the numerator, reflecting the similarity between an anchor point and its positive counterpart. The key distinction arises in the denominator, particularly in the treatment and summation of the negative samples. In InfoNCE, the negatives are typically all other samples that do not match the anchor’s positive pair. In contrast, our Smooth InfoNCE Loss introduces a $\lambda\leq 1$ parameter that effectively smoothens the contribution of each negative sample in the batch, apart from the anchor’s match.
III-C3 Self-Distillation with No Labels (DINO)
@@
III-D Backbone Neural Networks
@@ IV Experimental Results
In these metrics, TN (True Negative) and TP (True Positive) refer to the number of correctly classified negative and positive cases, respectively. Specifically, TN denotes cases that are correctly identified as negative, and TP indicates cases accurately recognized as positive. On the other hand, FP (False Positive) and FN (False Negative) represent the instances of incorrect predictions. FP occurs when a negative case is incorrectly predicted as positive, while FN happens when a positive case is wrongly classified as negative. These four elements are pivotal in assessing the model’s ability to accurately distinguish between positive and negative cases.
@@
TABLE II: A comparison of self-supervised learning with different backbone
@@
The table II comprehensively evaluates various machine learning models across different learning paradigms and data proportions. It compares the performance of MLP, FCNN, BiLSTM, and Transformer models, each trained under fully supervised, masking, contrastive, and DINO self-supervised learning paradigms. The performance metrics include Accuracy (Acc), Precision (Pre), Recall (Rec), and F1 Score (F1), evaluated at four different ratios of annotated data: 2.5%, 5%, 7.5%, and 10%. Table II shows that the Transformer model, particularly when trained with SSL approaches like masking, contrastive, and DINO, consistently achieves high scores across all metrics, indicating its robustness and adaptability in learning from annotated and unannotated data. It significantly improves the Transformer’s performance compared to supervised learning cases. For example, under the DINO self-supervised paradigm, the Transformer model shows remarkable performance with an accuracy and F1 score of 0.97 and 0.91, respectively, at the 2.5% data ratio and maintains similar high performance across other data ratios. In contrast, while MLP and BiLSTM models perform best under fully supervised learning, their performance does not significantly improve with SSL techniques. The FCNN model, however, shows some improvement with SSL compared to fully supervised learning, but not to the extent seen with the Transformer model. Overall, these results underline the effectiveness of SSL, particularly with the Transformer model, in handling various proportions of annotated and unannotated data, outperforming traditional supervised methods and other neural network architectures like MLP, BiLSTM, and FCNN in terms of all evaluation metrics.
@@
TABLE III: Transformer with self-supervised learning regarding different frameworks
@@
Table III showcases a comparative analysis of the Transformer model’s efficacy when trained under different SSL frameworks—masking, contrastive, and DINO—relative to traditional fully supervised training. At the outset, with only 2.5% annotated data, the SSL models generally surpass the fully supervised model on all counts. The contrastive and DINO approaches, in particular, demonstrate marked improvements, with the contrastive framework achieving top scores for accuracy and F1. As we escalate the annotated data to 5% and 7.5%, the contrastive learning framework notably maintains its high scores, aligning with the DINO approach in terms of accuracy and F1, and consistently outshining the fully supervised model. Upon reaching 10% annotated data, this pattern persists, with SSL frameworks, especially the contrastive and DINO, proving more efficient than fully supervised learning. They register high F1, indicative of their capacity to effectively harness annotated and unannotated data.
@@
TABLE IV: A comparison of contrastive loss from contrastive learning with Transformer
@@
Table IV evaluates the Transformer model using contrastive loss functions, including NT-Xent, SwCE, InfoNCE, and Smooth InfoNCE, revealing that with just 2.5% annotated data, NT-Xent and Smooth InfoNCE excel in accuracy and F1 score. As the annotated portion increases to 5% and 7.5%, these losses, along with SwCE, deliver top F1 scores, demonstrating a balanced precision-recall trade-off. At 10% annotation, they continue to showcase high accuracy and precision, with Smooth InfoNCE maintaining a notable F1 score. Synthesizing these insights, it is evident that contrastive loss functions significantly enhance the Transformer model’s performance, especially when annotated data is scarce. NT-Xent and Smooth InfoNCE, in particular, demonstrate consistent strength across all metrics, underlining their effectiveness within the Transformer’s contrastive learning framework.
Complementing Table IV, Figure 4 graphically shows the performance of the Transformer model under various contrastive losses. At the 2.5% annotation level, NT-Xent and Smooth InfoNCE manifest as superior performers. As the proportion of annotated data increases, these two losses retain their prominence, suggesting their efficacy in contexts with limited annotations. By the time annotated data reaches 10%, performances across the different loss functions begin to blend, with Smooth InfoNCE maintaining its strong precision, which could indicate a particular proficiency in generating confident predictions. Those results highlight the efficacy of contrastive loss functions in training Transformers, particularly in data-constrained environments. The consistently high performance of Smooth InfoNCE and NT-Xent suggests these methods are especially potent, showcasing the transformative potential of contrastive learning to boost Transformer models when resources are limited.
@@
Furthermore, motivated by adapting the unsupervised learning from Azar et al. [45], and our previous study [46], we continue to experiment with Transformer for different learning paradigms, including supervised, unsupervised, and self-supervised contrastive learning (with Smooth InfoNCE loss). Table V, and Fig. 6 present a comparative analysis of a Transformer model’s performance across three learning paradigms: supervised, unsupervised (using Autoencoders, AE), and self-supervised. Across all metrics, self-supervised learning exhibits consistently high performance, particularly excelling over supervised learning. Even with a minimal 2.5% of annotated data, self-supervised learning matches the top accuracy and precision shown by unsupervised learning while significantly surpassing supervised learning, especially in recall and F1 scores. As the amount of annotated data increases, the self-supervised Transformer maintains its superiority over supervised learning and demonstrates marked improvements over the unsupervised AE approach. This trend is consistent with up to 10% annotated data, where self-supervised learning still shows the highest metrics, confirming its effectiveness in leveraging both labeled and unlabeled data to improve performance, particularly in scenarios with limited annotations.
@@
TABLE V: A comparison of Transformer’s performance with different learning paradigm

### v2

@@ III-D1 Masking
Then, the SSL framework for data reconstruction from the masked data utilizes the masked data generated by the previous pseudo-algorithm. This framework aims to train a model to predict the original data from its masked version, thereby learning a robust representation of the data. Through this procedure, Algorithm 2, the model learns to predict the missing values in the masked data by reconstructing the original data. As the training progresses, the model becomes better at understanding the underlying patterns and structures in the data, even when some information is obscured. This SSL training paradigm effectively leverages the data as the supervisory signal, bypassing the need for external labels and allowing the model to learn unsupervised. This is particularly powerful for utilizing large unlabeled datasets to train models for tasks where labeled data is scarce to obtain.
III-D2 Contrastive Learning
@@
Contrastive learning is an SSL technique that teaches a model to distinguish between similar and dissimilar data points [30]. By doing so, the model learns rich, discriminative data representations without needing explicit labels. The essence of contrastive learning lies in its loss functions Eq. 6, 7, and 8, which drive the model to minimize the distance between positive pairs (similar items) and maximize the distance between negative pairs (dissimilar items). Here are three contrastive loss functions that are commonly used. Let’s denote $L_{ij}$ represents the loss for a positive pair of examples $(i,j)$. $\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)$ is the exponential function applied to the similarity score between the encoded representations $z_{i}$ and $z_{j}$. $\tau$ is the temperature parameter that scales the similarity score and controls the separation of the distribution of positive and negative examples.
@@
Normalized Temperature-scaled Cross Entropy (NT-Xent)[31]: This loss function is central to many contrastive learning algorithms and has been popularized by its use in SimCLR. It normalizes feature vectors and scales the dot product between them with a temperature parameter, encouraging the model to identify positive pairs among negative samples.
$L_{NT-Xent}=-\log\left(\frac{\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)}{\sum^{2N}_{k=1}1_{[k\neq i]}\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{k}\right)/\tau\right)}\right)$ (6)
NT-Xent scales the similarity scores and, by lowering the temperature, encourages the model to focus more on ”hard” negatives (i.e., those with high similarity to the anchor).
  • Uses the entire batch as potential negatives, making it very effective with large batch sizes.
  • The indicator function $1_{[k\neq i]}$ ensures that the similarity of the anchor with itself is excluded from the denominator.
  • The temperature parameter $\tau$ is used to control the separation between positives and negatives.
TABLE I: Comparison of Contrastive Loss Functions
Loss Function Focus Denominator (Negatives) Advantages Disadvantages
NT-Xent Identifying positive pairs in the presence of many negatives All other samples in the batch, excluding itself. Effective with large batch sizes. Sensitive to batch size and temperature parameters.
InfoNCE Learning representations by maximizing mutual information Explicit set of negative samples. Strong connection to mutual information estimation. Requires careful negative sampling.
SwCE Matching positive embedding to negatives Embeddings are swapped; matches positive to other negatives. Symmetric treatment of views can enhance robustness. Complexity in the formulation and swapping mechanism.
@@
Information Noise-Contrastive Estimation (InfoNCE) [32, 33]: InfoNCE is a variant that focuses on distinguishing a positive pair from a set of negatives. It’s designed to learn energy-based models and has been effectively used in representation learning, improving the quality of learned representations.
$L_{InfoNCE}=-\log\left(\frac{\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)}{\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)+\sum\limits_{\begin{subarray}{c}k=1\end{subarray}}^{N}\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{k}^{neg}\right)/\tau\right)}\right)$ (7)
InfoNCE encourages high mutual information between the anchor and its positive pair, maximizing the similarity for the positive while minimizing similarity with negative samples.
  • Focuses explicitly on the positive pair and contrasts it against a predefined set of negatives.
  • Each positive pair competes against $N$ negatives, typically resulting in better feature separation in high-dimensional spaces.
  • The denominator sums over the positives and negatives, which ensures that the loss is bounded and smoothly differentiable.
@@
Swapped Cross Entropy (SwCE)[34]: Introduced in SwAV, this function enforces consistency between cluster assignments produced by a pair of encoded views of the same image. It uses a swapping mechanism that leads to more robust representations by leveraging the invariance in the data.
$L_{SwCE}=-\log\left(\frac{\exp\left(\text{sim}\left(\mathbf{z}_{i},\mathbf{z}_{j}\right)/\tau\right)}{\sum^{N}_{k=1}\exp\left(\text{sim}\left(\mathbf{z}_{j},\mathbf{z}_{k}\right)/\tau\right)}\right)$ (8)
SwCE creates a symmetry between the views, encouraging the model to treat both views equally. This swapping mechanism can lead to more robust and balanced embeddings.
  • It flips the anchor-positive relationship, leading to a more balanced loss.
  • The denominator sums over the similarities of the positive embedding ($z_{j}$) to all other negative embeddings.
These loss functions are integral to the SSL framework, enabling models to effectively leverage large amounts of unlabeled data to learn useful data representations. Each loss function has advantages and can be chosen based on the specific requirements summarized in the Table. I.
Within this framework, we introduce Smooth InfoNCE, an extension of the standard InfoNCE loss that aims to tackle the challenges of overfitting and sensitivity to negative samples, particularly in limited data scenarios. While the traditional InfoNCE loss treats all negative samples with equal weight, this can lead to potential overfitting and an overly confident separation between positive and negative pairs, especially when the set of negative samples is small or lacks diversity. This behavior risks the model learning to ”memorize” negatives rather than capturing meaningful representations. The motivation behind the Smooth InfoNCE (Eq. 9), is as follows:
  • Overfitting to Negatives in Limited Data Scenarios: In standard InfoNCE, the contribution of negative samples is uniform. In cases of limited data, there is a risk of the model overfitting to these negatives. If the same small set of negatives is repeatedly used, the model might become overly confident in separating positive and negative pairs. This results in the model memorizing the negatives rather than learning meaningful representations.
  • Smoothing Factor to Mitigate Overconfidence: The smoothing factor $\lambda$ in the denominator helps modulate the contribution of negative samples, thus preventing the model from becoming overly confident in its predictions. This is especially useful when the negative samples are few or noisy.
  • Generalization in Small Datasets: By tuning $\lambda$, we can ensure that the loss does not emphasize the negatives, enhancing the generalization and robustness of the learned features.
$L_{SInfoNCE}=-\log\left(\frac{\exp\left(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j})/\tau\right)}{\exp\left(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j})/\tau\right)+\lambda\sum\limits_{\begin{subarray}{c}k=1,k\neq i\end{subarray}}^{N}\exp\left(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{k})/\tau\right)}\right)$ (9)
Smooth InfoNCE Loss can be seen as a special case of the standard InfoNCE loss. Both losses share a commonality in the numerator, reflecting the similarity between an anchor point and its positive counterpart. The primary distinction between the Smooth InfoNCE loss and the standard InfoNCE loss lies in how negative samples are treated. In the standard InfoNCE loss, the denominator sums over all negative samples with a uniform contribution, which means that each negative example is given equal importance during training. This can be problematic in limited data scenarios or when the quality of negative samples varies, as it may lead to overemphasizing specific negatives and potentially cause overfitting.
In contrast, the Smooth InfoNCE loss introduces a smoothing factor $\lambda$ that scales the contribution of each negative sample. When $\lambda\leq 1$, the influence of negatives is effectively reduced, thereby preventing the model from placing excessive confidence in distinguishing positives from these specific negative examples. This change acts as a form of regularization, making the loss function more resilient to noisy or ambiguous negative samples. As a result, the model learns more balanced representations, mitigating overfitting and enhancing generalization, especially in scenarios where the set of negative samples is small, noisy, or less diverse.
III-D3 Self-Distillation with No Labels (DINO)
@@
III-E Backbone Neural Networks
@@ IV Experimental Results
In these metrics, TN (True Negative) and TP (True Positive) refer to the number of correctly classified negative and positive cases, respectively. Specifically, TN denotes cases that are correctly identified as negative, and TP indicates cases accurately recognized as positive. On the other hand, FP (False Positive) and FN (False Negative) represent the instances of incorrect predictions. FP occurs when a negative case is incorrectly predicted as positive, while FN happens when a positive case is wrongly classified as negative. These four elements are pivotal in assessing the model’s ability to distinguish between positive and negative cases accurately.
TABLE II: Parameters setting of classifiers
Hyperparameters Transformer LSTM FCNN MLP
Hidden layers 4 2 3 3
Number of neurons 128 500 64 500
Number of multi-heads attention 4 N/A N/A N/A
Batch size 96 96 96 96
Dropout 0.25 0.3 0.25 0.3
Learning rate 6e-04 1e-04 1e-04 1e-04
Optimizer Adam Adam Adam Adam
Table II summarizes the hyperparameters used for different ML classifiers. The Transformer has 4 hidden layers with 128 neurons each, 4 multi-head attention mechanisms, a batch size of 96, a dropout rate of 0.25, a learning rate of 6e-04, and uses the Adam optimizer. The LSTM has two hidden layers with 500 neurons each, no multi-head attention, and a learning rate 1e-04. The FCNN and MLP classifiers share settings similar to those of the LSTM, including batch size, dropout rate, learning rate, and optimizer. Consequently, one of the importance of hyperparameters tuning of InfoNCE should be noted as follows:
  • Temperature Parameter ($\tau$): The temperature controls the sharpness of the softmax distribution. A low value (e.g., 0.1) makes the distribution sharper and emphasizes the most similar pairs more, while a higher value smooths the distribution.
  • Smoothing Factor ($\lambda$): The $\lambda$ value should be selected based on the dataset size and nature of the negative samples. Start with $\lambda=0.5$ and increase/decrease it to observe the effect. A lower $\lambda$ will reduce the impact of negatives, which is helpful if the negatives are noisy or very similar to positives.
@@
TABLE III: A comparison of self-supervised learning with different backbone
@@
The table III comprehensively evaluates various machine learning models across different learning paradigms and data proportions. It compares the performance of MLP, FCNN, BiLSTM, and Transformer models, each trained under fully supervised, masking, contrastive, and DINO self-supervised learning paradigms. The performance metrics include Accuracy (Acc), Precision (Pre), Recall (Rec), and F1 Score (F1), evaluated at four different ratios of annotated data: 2.5%, 5%, 7.5%, and 10%. Table III shows that the Transformer model, particularly when trained with SSL approaches like masking, contrastive, and DINO, consistently achieves high scores across all metrics, indicating its robustness and adaptability in learning from annotated and unannotated data. It significantly improves the Transformer’s performance compared to supervised learning cases. For example, under the DINO self-supervised paradigm, the Transformer model shows remarkable performance with an accuracy and F1 score of 0.97 and 0.91, respectively, at the 2.5% data ratio and maintains similar high performance across other data ratios. In contrast, while MLP and BiLSTM models perform best under fully supervised learning, their performance does not significantly improve with SSL techniques. The FCNN model, however, shows some improvement with SSL compared to fully supervised learning, but not to the extent seen with the Transformer model. Overall, these results underline the effectiveness of SSL, particularly with the Transformer model, in handling various proportions of annotated and unannotated data, outperforming traditional supervised methods and other neural network architectures like MLP, BiLSTM, and FCNN in terms of all evaluation metrics.
@@
TABLE IV: Transformer with self-supervised learning regarding different frameworks
@@
Table IV showcases a comparative analysis of the Transformer model’s efficacy when trained under different SSL frameworks—masking, contrastive, and DINO—relative to traditional fully supervised training. At the outset, with only 2.5% annotated data, the SSL models generally surpass the fully supervised model on all counts. The contrastive and DINO approaches, in particular, demonstrate marked improvements, with the contrastive framework achieving top scores for accuracy and F1. As we escalate the annotated data to 5% and 7.5%, the contrastive learning framework notably maintains its high scores, aligning with the DINO approach in terms of accuracy and F1, and consistently outshining the fully supervised model. Upon reaching 10% annotated data, this pattern persists, with SSL frameworks, especially the contrastive and DINO, proving more efficient than fully supervised learning. They register high F1, indicative of their capacity to effectively harness annotated and unannotated data.
@@
TABLE V: A comparison of contrastive loss from contrastive learning with Transformer
@@
Table V evaluates the Transformer model using contrastive loss functions, including NT-Xent, SwCE, InfoNCE, and Smooth InfoNCE, revealing that with just 2.5% annotated data, NT-Xent and Smooth InfoNCE excel in accuracy and F1 score. As the annotated portion increases to 5% and 7.5%, these losses, along with SwCE, deliver top F1 scores, demonstrating a balanced precision-recall trade-off. At 10% annotation, they continue to showcase high accuracy and precision, with Smooth InfoNCE maintaining a notable F1 score. Synthesizing these insights, it is evident that contrastive loss functions significantly enhance the Transformer model’s performance, especially when annotated data is scarce. NT-Xent and Smooth InfoNCE, in particular, demonstrate consistent strength across all metrics, underlining their effectiveness within the Transformer’s contrastive learning framework.
Complementing Table V, Figure 4 graphically shows the performance of the Transformer model under various contrastive losses. At the 2.5% annotation level, NT-Xent and Smooth InfoNCE manifest as superior performers. As the proportion of annotated data increases, these two losses retain their prominence, suggesting their efficacy in contexts with limited annotations. By the time annotated data reaches 10%, performances across the different loss functions begin to blend, with Smooth InfoNCE maintaining its strong precision, which could indicate a particular proficiency in generating confident predictions. Those results highlight the efficacy of contrastive loss functions in training Transformers, particularly in data-constrained environments. The consistently high performance of Smooth InfoNCE and NT-Xent suggests these methods are especially potent, showcasing the transformative potential of contrastive learning to boost Transformer models when resources are limited.
@@
Furthermore, motivated by adapting the unsupervised learning from Azar et al. [45], and our previous study [46], we continue to experiment with Transformer for different learning paradigms, including supervised, unsupervised, and self-supervised contrastive learning (with Smooth InfoNCE loss). Table VI, and Fig. 6 present a comparative analysis of a Transformer model’s performance across three learning paradigms: supervised, unsupervised (using Autoencoders, AE), and self-supervised. Across all metrics, self-supervised learning exhibits consistently high performance, particularly excelling over supervised learning. Even with a minimal 2.5% of annotated data, self-supervised learning matches the top accuracy and precision shown by unsupervised learning while significantly surpassing supervised learning, especially in recall and F1 scores. As the amount of annotated data increases, the self-supervised Transformer maintains its superiority over supervised learning and demonstrates marked improvements over the unsupervised AE approach. This trend is consistent with up to 10% annotated data, where self-supervised learning still shows the highest metrics, confirming its effectiveness in leveraging both labeled and unlabeled data to improve performance, particularly in scenarios with limited annotations.
@@
TABLE VI: A comparison of Transformer’s performance with different learning paradigm